# thanks: This paper has been submitted in conjunction with its companion Algeri et al. [1]

### v1

@@ Appendix B The covariance of the angular power spectrum
$\begin{split}&|(\hat{\Sigma}_{{f_{b}},s})_{\ell,\ell^{\prime}}-({\Sigma}_{{f_{b}},s})_{\ell,\ell^{\prime}}|\\ \leq&\sum_{m,m^{\prime}}\frac{2\,\left|({a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}})(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\right|}{(1+2\ell)(1+2\ell^{\prime})}\\ \leq&{C}\cdot\|\Gamma_{{f_{b}},s}^{-1}\|_{op}\sum_{m,m^{\prime}}|{a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}|\\ \leq&{C}\cdot\zeta\sum_{m,m^{\prime}}|{a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}|.\end{split}$
Since for any ${f_{b}}$ and any $|m|\leq\ell\leq\ell_{max}$
$\hat{a}_{{f_{b}},\ell m}\;\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\;a_{{f_{b}},\ell m}\quad\text{as }S\to\infty,$
@@
$\quad\ \langle U_{\bm{a},\bm{b}}\bm{x},U_{\bm{a},\bm{b}}\bm{y}\rangle$
@@
$=\langle\bm{x},\bm{y}\rangle-\frac{\langle\bm{a}-\bm{b},\bm{x}\rangle\langle\bm{a}-\bm{b},\bm{y}\rangle}{(1-\langle\bm{a},\bm{b}\rangle)^{2}}\left(||\bm{a}-\bm{b}||^{2}-2-2\langle\bm{a},\bm{b}\rangle\right)$

### v3

@@ Appendix B The covariance of the angular power spectrum
$\begin{split}&|(\hat{\Sigma}_{{f_{b}},s})_{\ell,\ell^{\prime}}-({\Sigma}_{{f_{b}},s})_{\ell,\ell^{\prime}}|\\ \leq&\sum_{m,m^{\prime}}\frac{2\,\left|({a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}})(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\right|}{(1+2\ell)(1+2\ell^{\prime})}\\ \leq&{C}\cdot\|\Gamma_{{f_{b}},s}^{-1}\|_{\rm op}\sum_{m,m^{\prime}}|{a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}|\\ \leq&{C}\cdot\zeta\sum_{m,m^{\prime}}|{a}^{*}_{{f_{b}},\ell m}{a}_{{f_{b}},\ell^{\prime}m^{\prime}}-\hat{a}^{*}_{{f_{b}},\ell m}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}|.\end{split}$ (132)
Since for any $f_{b}$} and any $|m|\leq\ell\leq\ell_{\rm max}$
$\hat{a}_{{f_{b}},\ell m}\;\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\;a_{{f_{b}},\ell m}\quad\text{as }S\to\infty,$ (133)
@@
$\quad\ \langle U_{\bm{a},\bm{b}}\bm{x},U_{\bm{a},\bm{b}}\bm{y}\rangle$ (138)
@@
[…]